# Look Before you Leap: Estimating LLM Benchmark Scores from Descriptions

### v1

@@ 5 Experiments
We compare against three families of baselines. First, we report the overall test-split mean across all metrics (Section 5.1) and the per-metric test-split means (Section 5.2). Because no training or validation set is available, the test-set average serves as a strong baseline for comparison. Second, we evaluate text-encoder–based regressors. We embed each description and train classical models—k-nearest neighbors (Guo et al., 2003) and XGBoost (Chen & Guestrin, 2016)—on these embeddings. With no dedicated training set, we use k-fold cross-validation (Fushiki, 2011): fit on k-1 folds, evaluate on the held-out fold, and average over folds. For embeddings, we use E5-Mistral-7B (Wang et al., 2023), a leading open-source encoder.
@@
We evaluate both proprietary and open-source models in reasoning and non-reasoning settings to isolate the contribution of latent chain-of-thought style reasoning from raw generation quality. On the proprietary side we use GPT-5 (OpenAI, 2025) as a reasoning model. On the open-source side we include reasoning-enabled model Qwen3-32B (Yang et al., 2025), alongside non-reasoning models Qwen2.5-32B, Qwen2.5-72B (Team, 2024). Using matched prompts and decoding settings across these pairs lets us attribute gains specifically to reasoning capabilities. We prompt the model to output its final estimate in boxed{} with a brief rationale, then parse the boxed value during post-processing. Predictions are produced via greedy decoding with default generation settings.
Search Retrieval
All models share a single custom retriever over the arXiv API [^3] https://info.arxiv.org/help/api/index.html that returns canonical arXiv PDF links. To prevent leakage, we exclude each record’s source papers using identifier during retrieval. We initially tested GPT-5’s built-in web search—more token-efficient but corpus-uncontrolled—which frequently surfaced source papers; in an audit, $\sim\!70\%$ of evaluation dataset-source papers became identifiable, undermining realism. Consequently, despite higher processing cost, we rely on the custom retriever for GPT-5. We report results with retrieval on and off to assess gains from literature grounding beyond the model’s prior. Each query returns at most two documents, and the agent may issue up to four retrievals under ReAct-style iterative prompting (Yao et al., 2023), implemented via Qwen-Agent  [^4] https://github.com/QwenLM/Qwen-Agent and OpenAI Response API  [^5] https://platform.openai.com/docs/api-reference/responses.
@@ 5.1 Evaluation Results
Model MAE $\mathbf{r}$
@@
Test set average 21.4
E5-Mistral + kNN 19.9 0.30
E5-Mistral + XGBoost 20.3 0.30
@@
GPT-5 14.7 0.69
GPT-5 + Search 14.0 0.73
Qwen3-32B 19.6 0.40
Qwen3-32B + Search 20.8 0.31
Qwen2.5-32B 23.5 0.40
Qwen2.5-32B + Search 22.0 0.24
Qwen2.5-72B 22.4 0.44
Qwen2.5-72B + Search 21.4 0.34
@@
Table 3: Aggregated results (767 instances); Search = arXiv retrieval. indicates significance at $p<0.05$ (one-sided binomial sign test).
We report overall results in Table 3. Except for GPT-5 (OpenAI, 2025), no baseline meaningfully outperforms the test-set global mean. GPT-5 with Search achieves the best overall MAE (14.0) yet still leaves substantial headroom, while GPT-5 without Search also performs well, suggesting that pretrained knowledge and difficulty cues already carry useful signal. Classical text-encoder baselines—kNN and XGBoost with k-fold CV—do not meaningfully outperform the global mean. This suggests that effective forecasting most likely requires (i) reasoning, (ii) representations that capture fine-grained cues in long descriptions, and (iii) knowledge beyond simple interpolation from seed results. Among Qwen variants (Team, 2024; Yang et al., 2025), Qwen3-32B (reasoning) leads, however enabling Search gives inconsistent MAE and can degrade it, indicating brittle tool use. Qwen2.5-72B outperforms Qwen2.5-32B but still trails Qwen3-32B (reasoning).
5.2 Further Analysis Across Data Subsets
@@
We report per-metric MAE against a test-calibrated per-metric mean baseline—stronger than a single global mean—and compare GPT-5 (with/without Search) across seven metric groups, given its accurate, non-trivial forecasts. Figure 3 reports per-metric MAE. GPT-5 generally outperforms the metric-calibrated test-set mean baseline. The Accuracy subset—our largest slice—mirrors the overall trend. By contrast, for BLEU and ROUGE, whose gold scores have low variance, LLMs sometimes fall below the baseline, reflecting the difficulty of forecasting fine-grained metric values.
Peer-reviewed arXiv
# Samples 520 247
GPT-5 14.6 14.5
GPT-5 + Search 13.8 14.4
Table 4: Peer-reviewed vs. arXiv: sample counts and MAE.
Figure 3: Per-metric MAE (↓) by model, grouped by evaluation metric. The red bar marks the per-metric average with its absolute MAE label; other bars show delta to that average. The test set comprises 767 instances: 419 Accuracy, 152 F1, 56 Recall, 42 Precision, 35 EM, 32 ROUGE, and 31 BLEU.
We further stratify results by venue to assess whether publication status correlates with predictability. We compare two groups: Peer-reviewed samples appearing in refereed venues, arXiv postings in CoRR preprints. As shown in Table 4, the peer-reviewed subset and arXiv subset exhibit very similar MAE for GPT-5. Adding Search gives modest gains on the peer-reviewed subset vs. arXiv, hinting at answer leakage; accordingly, we also evaluate a streaming setting (Section 7) that forecasts on newly released datasets before citations/indexing. Overall differences are small, suggesting venue has limited effect and task characteristics dominate.
@@ 5.3 Human prediction results
In the previous sections, we demonstrated that LLMs can generate accurate predictions from dataset descriptions. A natural next question is how these predictions compare to those made by human experts. To explore this, we asked six graduate students and one professor of computer science, all specializing in NLP and machine learning, to each predict 20 accuracies based solely on the same text descriptions provided to models. On this 140-item subset, human predictions correlated moderately with ground truth ($r=0.403$) and outperformed the mean baseline in terms of MAE (19.6 vs. 21.7). Nevertheless, GPT-5 with search achieved substantially stronger results, indicating that while humans can do the task reasonably well, advanced reasoning models achieve markedly lower MAE. Although human forecasters might improve on this task with additional training (Tetlock et al., 2014), the fact that LLMs already outperform researchers on the performance prediction task suggests benchmark performance prediction models may provide meaningful value to practitioners when planning experiments.
6 Confidence and Retrieval Analysis
Figure 4: MAE (left axis) and coverage (right axis) as a function of confidence thresholds ($\geq$ band lower bound) for GPT-5 and Qwen3-32B, each with/without search. Higher thresholds retain fewer examples (lower coverage) and generally correspond to lower MAE.
Some experimental outcomes may be inherently difficult to predict, raising the question of whether models can reliably estimate their own prediction accuracy. To examine this, we elicit a scalar, self-reported confidence for each prediction (following Yoon et al. (2025)) and evaluate calibration by filtering to items with confidence $\geq\tau$. For each $\tau$, we report the MAE on the retained subset together with its coverage. This gives a simple view of the error–coverage trade-off. We report confidence curves for the accuracy subset, which dominates Precog.
Figure 4 shows results on the accuracy subset for GPT-5 and Qwen3-32B (with and without search). As the confidence threshold increases, coverage declines and MAE tends to decrease, indicating that higher reported confidence is, on average, associated with more accurate forecasts. Differences between “with search” and “without search” are present but modest and may vary by threshold. Overall, these curves suggest reasonable alignment between self-reported confidence and error without making strong claims about perfect calibration.
GPT5 Qwen3
@@ 6 Confidence and Retrieval Analysis
Table 6: Search-tool usage and token diversity; tokens counts measured with the GPT-4o tokenizer. (all) aggregates all formulated queries; (random) uses one random query per instance.
6.1 Retrieval Behavior
We analyze retrieval behavior via tool-call counts and query token diversity (tokenized with GPT-4o (Hurst et al., 2024)). As summarized in Table 6, GPT-5 issues more—and more variable—search calls per example than Qwen3-32B. GPT-5’s queries also show higher lexical diversity versus Qwen. These descriptive differences suggest GPT-5 tends to aggregate evidence across more—and more varied—queries, whereas Qwen’s lighter tool use may correspond to more limited gains.
@@ 7 Streaming Prediction
Using Precog with a leakage-mitigating custom retriever, GPT-5 outperforms the per-metric test-set mean baseline. Nevertheless, residual answer leakage cannot be entirely excluded—for instance, when subsequent papers copy baseline results from the original publication, or when models memorize results during pre- or post-training. To further validate our experimental setup under more realistic conditions, we adopt a streaming prediction setting, which tests models on their ability to predict performance on newly reported arXiv results immediately upon release, before they can be indexed by search engines or incorporated into training.
@@
Figure 5: Streaming prediction error boxplots.
We collect candidates from arXiv’s official interfaces (API/RSS) for cs.CL, cs.LG, and cs.AI, and fetch PDFs via the recommended hosts, respecting robots.txt and published rate limits; we also notified arXiv staff in advance (see Appendix F). [^6] Recent lists: https://arxiv.org/list/cs.CL/recent, https://arxiv.org/list/cs.LG/recent, https://arxiv.org/list/cs.AI/recent.
A daily, low-impact pipeline then (i) identifies papers containing results that match our schema, (ii) extracts table records, (iii) builds redacted, schema-aligned descriptions (at most two per record), and (iv) issues predictions—end-to-end before third-party search engines index the papers. In this setting, we restrict to records whose metrics map to our seven metric groups (normalized to 0–100) and expand the model axis to the GPT family (beyond GPT-4/4o used on Precog). We permit GPT-5’s built-in search here, as leakage from prior literature is negligible for freshly released papers. We collected 52 papers containing 79 previously unseen experimental records and datasets between 2025-09-05 and 2025-09-23.
Results
@@ Conclusion.
We recast evaluation as forecasting and introduce Precog, a leakage-mitigating corpus of redacted, schema-aligned description–performance pairs. Search-augmented reasoning with GPT-5 produces accurate forecasts with calibrated uncertainty, and remains effective in a streaming setting where results are predicted for newly released papers prior to broad indexing. These findings indicate that anticipatory evaluation is feasible and practically useful for estimating task difficulty, prioritizing experiments, and reducing compute.
@@
Despite progress, open-source baselines lag behind GPT-5 in both accuracy and calibration. They show brittle tool use (inconsistent gains from retrieval), weaker representations of fine-grained cues in long descriptions, and higher variance across metric families (e.g., small-variance metrics like BLEU/ROUGE). Improved query planning, evidence selection, and uncertainty modeling are needed for reliable deployment.
Discussion: feasibility and utility.
[…]
@@
To accelerate scientific discovery, we advocate purpose-built regressors that leverage LLM reasoning. In particular, reasoning-aware regression trains a regression model on top of reasonig LLMs, while better use of retreival tool focuses on learned query planning and selective evidence use. Developing these components on top of Precog and the streaming benchmark should yield more accurate, trustworthy, and cost-efficient forecasting systems.

### v2

@@ 5 Experiments
We compare against three families of baselines. First, we report the overall test-split mean across all metrics (Section 5.1) and the per-metric test-split means (Section 5.2). Because no training or validation set is available, the test-set average serves as a strong baseline for comparison. Second, we evaluate text-encoder–based regressors. We embed each description and train classical models—k-nearest neighbors (guo2003knn) and XGBoost (chen2016xgboost)—on these embeddings. With no dedicated training set, we use k-fold cross-validation (fushiki2011estimation): fit on k-1 folds, evaluate on the held-out fold, and average over folds. We set k = 10 and used E5-Mistral-7B (wang2023improving), which is a strong dense embedding encoder baseline. Hyperparameters were tuned on a held-out validation split from the 2023–2024 version of Precog.
@@
We evaluate proprietary and open-source LLMs in both reasoning and non-reasoning modes to disentangle chain-of-thought effects from base generation quality. On the proprietary side, we use GPT-5 (openai2025gpt5) in its reasoning configuration; on the open-source side, we include reasoning-enabled Qwen3-32B (yang2025qwen3) and non-reasoning Qwen2.5-32B/72B (team2024qwen2). For reasoning-capable models, we also run a no-reasoning variant (GPT-5 with minimal reasoning; Qwen3 with no-thinking) to quantify the contribution of explicit reasoning. We match prompts and decoding settings across modes so that performance differences can be attributed to reasoning. In all settings, the model produces a brief rationale and encloses its final estimate in boxed{}, which we parse in post-processing. GPT-5 uses default decoding hyperparameters, and Qwen3 uses the original authors’ recommended settings.
@@ 5.1 Evaluation Results
2023–2024 2025
Model MAE $\downarrow$ $\mathbf{r}$ $\uparrow$ MAE $\downarrow$ $\mathbf{r}$ $\uparrow$
@@
Test set average 21.4 23.0
E5-Mistral + kNN 19.9 0.30 21.2 0.22
E5-Mistral + XGBoost 20.3 0.30 21.4 0.20
@@
GPT-5 14.7 0.69 14.6 0.62
GPT-5 + minimal-reasoning 15.0 0.68 15.5 0.52
Qwen3-32B 19.6 0.40 19.6 0.41
Qwen3-32B + no-thinking 22.9 0.37 22.5 0.37
Qwen2.5-32B 23.5 0.40 21.4 0.39
Qwen2.5-72B 22.4 0.44 19.5 0.43
@@
Table 3: Prediction performance on Precog for 2023–2024 (767) and 2025 (1,523 instances). indicates significance at $p<0.05$ (one-sided binomial sign test) for the corresponding column.
We summarize the main validation results in Table 3. With the exception of GPT-5 (openai2025gpt5), no baseline meaningfully outperforms the test-set global mean. GPT-5 achieves the lowest overall MAE (14.6), but still leaves substantial headroom. Performance is similar on the 2023–2024 and leakage-free 2025 splits, suggesting that prediction accuracy is not driven by training-data memorization. Moreover, reasoning ablations for GPT-5 and Qwen3 show that enabling reasoning mode improves performance prediction. Classical text-encoder baselines—kNN and XGBoost with k-fold CV—also fail to beat the global mean by a meaningful margin. Among Qwen variants (team2024qwen2; yang2025qwen3), Qwen3-32B in reasoning mode performs best, and is comparable to Qwen2.5-72B despite using roughly half as many parameters.
5.2 Prediction Heterogeneity by Task, Metric, Target, and Setup
@@
We evaluate GPT-5 across seven metric groups by reporting per-metric MAE relative to a test-calibrated per-metric mean baseline, which is stronger than a single global mean. As shown in Figure 3, GPT-5 typically outperforms this baseline on both the 2023–2024 and 2025 splits. The Accuracy subset—our largest slice—closely tracks the overall trend. In contrast, for Rouge and Precision, whose gold scores have low variance, LLMs sometimes fall below the baseline, underscoring the difficulty of forecasting such fine-grained metrics.
Figure 3: Per-metric MAE (↓) by model, grouped by evaluation metric and computed separately on the 2023–2024 and 2025 data subsets. The red and orange bars show the per-metric mean, labeled with its absolute MAE; the other bars show deviations from that mean within the corresponding year-split subset.
GPT-5 Qwen3
Category $n$ MAE $\downarrow$ $\mathbf{r}$ $\uparrow$ MAE $\downarrow$ $\mathbf{r}$ $\uparrow$
Knowledge 970 14.5 0.68 19.3 0.41
Reasoning 1100 14.8 0.67 19.5 0.37
Math 131 15.0 0.74 21.4 0.45
Coding 170 15.1 0.73 20.4 0.55
Safety 63 15.8 0.74 19.6 0.53
IF 382 15.5 0.66 20.7 0.40
Multilingual 60 12.0 0.78 17.9 0.60
(a) By benchmark category.
Predictor Test Set $n$ MAE $\downarrow$ $\mathbf{r}$ $\uparrow$
GPT-5 Peer-reviewed 486 15.4 0.69
GPT-5 Not peer-reviewed 1036 14.5 0.68
Qwen3 Peer-reviewed 487 20.7 0.40
Qwen3 Not peer-reviewed 1036 19.1 0.42
(b) Peer-reviewed vs. non–peer-reviewed.
Predictor Target Model $n$ MAE $\downarrow$ $\mathbf{r}$ $\uparrow$
GPT-5 GPT-4 406 13.9 0.73
GPT-5 GPT-4o 1116 15.2 0.66
Qwen3 GPT-4 407 18.1 0.48
Qwen3 GPT-4o 1116 20.2 0.40
(c) By target model (GPT-4 vs. GPT-4o).
Predictor Test Set $n$ MAE $\downarrow$ $\mathbf{r}$ $\uparrow$
GPT-5 Established datasets 973 14.7 0.69
GPT-5 New datasets 549 15.1 0.67
Qwen3 Established datasets 974 19.1 0.42
Qwen3 New datasets 549 20.7 0.40
(d) Established datasets vs. New datasets.
Table 4: Subset ablation analyses on Precog 2025. Each panel reports MAE and Pearson $\mathbf{r}$ for GPT-5 and Qwen3-32B under different subset. Peer-reviewed entries are identified by matching result papers in DBLP. We define established datasets records as results on previously introduced datasets, and new datasets records as results reported in papers that introduce a new dataset.
We analyze heterogeneity on the Precog 2025 snapshot along benchmark category, publication status, target model, and established datasets vs. new datasets in Table 4. We define established datasets records as instances where the result reporting paper differs from the dataset paper, and new datasets records as instances where they are the same (the result paper introduced the dataset). Across categories, both GPT-5 and Qwen3-32B show similar patterns: multilingual benchmarks are easiest to forecast (lowest MAE, highest $r$), while math and instruction-following are harder, likely reflecting more diverse task formulations and evaluation setups. Performance is also very similar on peer-reviewed and non–peer-reviewed benchmarks, with slightly better MAE on non–peer-reviewed entries, suggesting that venue and “maturity” have limited impact.
Conditioning on the target model, both predictors perform better on GPT-4 than GPT-4o, indicating that newer or more variable targets are modestly harder to forecast, plausibly because GPT-4 results were more exposed during LLM training. Finally, MAE and correlation are nearly identical for established datasets vs. new datasets records, implying that the predictors transfer well from established datasets to newly introduced, previously unseen datasets.
@@ 5.3 Human prediction results
In the previous sections, we demonstrated that LLMs can generate accurate predictions from dataset descriptions. A natural next question is how these predictions compare to those of human experts, who are known to struggle with probability judgments and decisions under uncertainty (kahneman1972subjective). To explore this, we asked six graduate students and one professor of computer science, all specializing in NLP and machine learning, to each predict 20 accuracies based solely on the same text descriptions provided to models. On this 140-item subset, human predictions correlated moderately with ground truth ($r=0.403$) and outperformed the mean baseline in terms of MAE (19.6 vs. 21.7). Nevertheless, GPT-5 achieved substantially stronger results, indicating that while humans can do the task reasonably well, advanced reasoning models achieve markedly lower MAE. Although the study has a small sample size, the results primarily serve to illustrate the inherent difficulty of the task. While human forecasters might improve with additional training (tetlock2014forecasting), the fact that LLMs already outperform researchers on this performance prediction task suggests that benchmark prediction models have the potential to provide meaningful support to practitioners when planning experiments.
6 Confidence Calibration
Figure 4: MAE (left axis) and coverage (right axis) as a function of confidence thresholds (scores $\geq$ band lower bound) for GPT-5 and Qwen3-32B across different year splits (2023-2024 vs 2025). Higher thresholds retain fewer examples (lower coverage) and generally yield lower MAE. Confidence values are derived from a 10-bin categorical output produced by the verbalized confidence calibration.
Some experimental outcomes may be inherently difficult to predict, raising the question of whether models can reliably estimate their own prediction accuracy. To examine this, we elicit a scalar, self-reported confidence for each prediction (following yoon2025reasoning) and evaluate calibration by filtering to items with confidence $\geq\tau$. For each $\tau$, we report the MAE on the retained subset together with its coverage. This gives a simple view of the error–coverage trade-off. We report confidence curves for the accuracy subset, which dominates Precog.
Figure 4 shows results on the accuracy subset for GPT-5 and Qwen3-32B on different year-subset split. As the confidence threshold increases, coverage declines and MAE tends to decrease, indicating that higher reported confidence is, on average, associated with more accurate forecasts. The 2023–2024 split tends to be better calibrated—MAE drops more sharply as the confidence threshold increases—which may reflect that the evaluated LLM has seen more similar examples from this period during training. Overall, these curves suggest reasonable alignment between self-reported confidence and error without making strong claims about perfect calibration.
GPT-5 Qwen3-32B
Prediction performance (MAE / $\mathbf{r}$)
No Search 14.7 / 0.69 19.6 / 0.40
With Search 14.0 / 0.73 20.8 / 0.31
Search-tool usage & token diversity
@@ 6 Confidence Calibration
Table 6: Search-tool usage, token diversity, and prediction performance for GPT-5 and Qwen3-32B. Token counts are measured with the GPT-4o tokenizer. (all) aggregates all formulated queries; (random) uses one random query per instance.
7 Retrieval Analysis
We assess how retrieval contributes to performance prediction using tool-call counts and query token diversity (tokenized with GPT-4o (hurst2024gpt)). All models share a custom retriever over the arXiv API, [^2] https://info.arxiv.org/help/api/index.html which returns canonical arXiv PDF links. To avoid leakage, we exclude each record’s source papers by identifier during retrieval. We initially tried GPT-5’s built-in web search—which is more token-efficient but corpus-uncontrolled—and found that it frequently surfaced source papers: in an audit, roughly 70% of evaluation source papers became identifiable, so we instead use the custom retriever for GPT-5 despite its higher cost. We report results with retrieval enabled and disabled to measure gains from literature grounding beyond the model’s prior. Each query returns at most two documents, and the agent may issue up to four retrievals under ReAct-style prompting (yao2023react).
As shown in Table 6, search improves GPT-5’s performance prediction but not Qwen3’s. GPT-5 issues more—and more variable—search calls per example than Qwen3-32B, and its queries are more lexically diverse, suggesting it aggregates evidence across a broader set of retrieved documents, whereas Qwen’s lighter tool use limits potential gains. Yet when we asked an LLM to rate retrieved documents on a 5-point relevance scale, most were only “moderately relevant”: the retriever often surfaces high-level related work rather than papers that add much beyond pretraining and our structured descriptions. This helps explain why sparse retrieval has limited impact and can even hurt Qwen3 in some cases. Overall, retrieval is not the main source of forecasting gains in Precog; most signal comes from reasoning over curated descriptions plus the model’s own knowledge.
@@ 8 Streaming Prediction
Figure 5: Streaming prediction error boxplots of GPT-5+Web-Search.
Using Precog, GPT-5 outperforms the per-metric test-set mean baseline. Nevertheless, residual answer leakage cannot be entirely excluded—for instance, when subsequent papers copy baseline results from the original publication. Moreover, we disabled built-in web search because it risks leakage, for example by retrieving the target papers during prediction. To further validate our experimental setup under more realistic conditions, we adopt a streaming prediction setting, which tests models on their ability to predict performance on newly reported arXiv results immediately upon release, before they can be indexed by search engines or incorporated into training.
@@
We collect candidates from arXiv’s official interfaces (API/RSS) for cs.CL, cs.LG, and cs.AI, and fetch PDFs via the recommended hosts, respecting robots.txt and published rate limits; we also notified arXiv staff in advance (see Appendix G).
A daily pipeline then (i) identifies papers of interest, (ii) extracts table records, (iii) builds redacted descriptions, and (iv) issues predictions—end-to-end before third-party search engines index the papers. In this setting, we restrict to records whose metrics map to our seven metric groups and expand the model axis to the GPT family. We permit built-in web-search here, as leakage from prior literature is negligible for freshly released papers. We collected 52 papers containing 79 previously unseen experimental records and datasets between 2025-09-05 and 2025-09-23. During search, GPT’s web tool may surface the target paper. We match arXiv IDs and verify content to detect any such access, and we exclude those predictions to prevent leakage.
@@ Conclusion.
We recast evaluation as forecasting and introduce Precog, a leakage-mitigating corpus of redacted, schema-aligned description–performance pairs. GPT-5 produces moderately accurate forecasts with calibrated uncertainty, and remains effective in a streaming setting where results are predicted for newly released papers prior to broad indexing. These findings indicate that anticipatory evaluation is feasible and practically useful for estimating task difficulty, prioritizing experiments, and reducing compute.
Discussion: scope and limitations.
[…]
Discussion: cost and latency.
Using GPT-5 (medium reasoning, no search) via the Batch API costs about $0.01$ per query and averages $\sim$132 seconds per prediction with high parallelism. For open-source baselines (e.g., Qwen), we run single-GPU inference on an A40 with a maximum generation length of 32,76832{,}768 tokens; a single prediction averages $\sim$48 seconds. While not directly comparable to end-to-end benchmark runs, description-based forecasting is markedly more cost- and time-efficient, making it well suited for early screening and study design.
@@
Despite progress, open-source baselines lag behind GPT-5 in both accuracy and calibration. They show brittle tool use (inconsistent gains from retrieval), weaker representations of fine-grained cues in long descriptions, and higher variance across metric families (e.g., small-variance metrics like BLEU/ROUGE). Improved reasoning and uncertainty modeling are needed for reliable deployment.
@@
To accelerate scientific discovery, we advocate purpose-built regressors that leverage LLM reasoning. In particular, reasoning-aware regression trains a regression model on top of reasoning LLMs, while better use of retrieval tools focuses on learned query planning and selective evidence use. An important future direction is to make these forecasting systems robust for new models with very few or no prior evaluations, for example by incorporating stronger priors, actively collecting a small set of seed evaluations, or using richer hybrid text–metadata representations. Developing these components on top of Precog and the streaming benchmark should yield more accurate, trustworthy, and cost-efficient forecasting systems.